# Eccentric and cool: a high-spectral resolution view of

### v1

@@ 5 Discussion and conclusions
The implications of the planet’s high eccentricity have been extensively discussed in previous studies (Maire et al. 2019; Dupuy et al. 2021; Balmer et al. 2025). Dynamical interactions with a massive inner companion could in principle reproduce the observed eccentricity, but the required mass would exceed SPHERE detection limits (Maire et al. 2019). An alternative scenario involves secular resonances with the massive circumstellar disk, which can excite planetary eccentricities up to $\sim$0.6 even for modest disk eccentricities (Li and Lai 2023). Kozai–Lidov oscillations induced by a distant third body, such as the stellar companion GJ 3305 AB, can also generate large eccentricities (Kozai 1962; Lidov 1962). However, the wide separation of GJ 3305 AB implies a Kozai–Lidov timescale longer than the age of the system, making this scenario unlikely (Montet et al. 2015), except in the case of a very high stellar companion eccentricity ($\gtrsim 0.9$; see Appendix F). Kozai–Lidov oscillations with an undetected low-mass outer companion remain possible; given SPHERE/IRDIS detection limits (Chomez et al. 2025), such a companion would likely be Jovian or sub-Jovian ($\lesssim 1$$2,M_{\mathrm{Jup}}$). Continued astrometric and radial velocity monitoring will further constrain this scenario (Lacour et al. 2021).
@@ Appendix B Considering whether the results of the fourth epoch are biased
To highlight these correlated structures, it is convenient to compute the Autocorrelation Function (ACF) of the CCF. Fig. 4 presents the individual ACFs of the 4 CCFs. The first three epochs are very consistent. The last epoch, however, exhibits prominent correlated structures at velocities multiples of $\sim$ 11 $\mathrm{km}\,\mathrm{s}^{-1}$, which is consistent with the previous analysis. It also shows that the fringes cause a decrease in the Full Width at Half-Maximum (FWHM) of the ACF, compared to the other epochs. It means that the fringes bias the FWHM of the CCF to lower values. Since we expect the FWHM of the CCF and the $v\sin i$ to be correlated (see Brady et al. 2023, for example), we can conclude that the fringes likely bias the $v\sin i$ estimate of the last epoch towards lower values. However, they likely do not bias the RV as much. Provided that the template correlates better with the data than the fringes do, the CCF still provides a correct measurement of the RV of the planet.
@@
Figure 3: Power-spectral density (PSD) of the data filtered from the stellar speckles estimate of our four epochs. The PSD was lowpass filtered in order to focus on its low-frequency content. A secondary axis is displayed at the top, which corresponds to an equivalent velocity. A peak is visible around a resolution of 25 000 for each epoch, which corresponds to $\sim$12 $\mathrm{km}\,\mathrm{s}^{-1}$. This peak is particularly prominent in the data from the last epoch, when the planet was closer to the host star.
Figure 4: Normalized Autocorrelation Functions (ACF) of the individual CCFs. The first three epochs show highly consistent ACFs, whereas the last epoch shows deeply correlated structures in the CCF at multiples of 11 km/s. The vertical dashed lines are plotted at multiples of $\pm$ 11 $\mathrm{km}\,\mathrm{s}^{-1}$ to highlight the correlated structures.
@@ Appendix C Method for computing the cross-correlation functions
To compute the cross-correlation functions, we use the formalism defined in Zucker (2003). The cross-correlation function is defined as:
@@
Figure 5: Individual CCF for the 4 epochs.
@@ Appendix D Orbit posterior distribution
Fig. 6 presents the posterior distribution of the orbital parameters of 51 Eri b, without our RV measurements (in red) and with our RV measurements (in blue). The inclusion of the HiRISE RV measurements brings significant improvements to the orbital parameters. It also breaks the degeneracy between particular values of $\omega$ and $\Omega$ typically arising with orbital fittings using only astrometric data points.
@@ Appendix E Contrast in reflected and emitted light
with $R$ the radius of the planet, $a$ its semi-major axis, $A$ its albedo and $f$ the phase function. We considered 3 different values for the phase function to show the impact of the phase of the planet. For $A$, we considered a value of 0.5, similar to that of Jupiter in visible (Karkoschka 1998). The values of $a$ and $R$ have been determined from the orbital and atmospheric fit, respectively.
@@ Appendix F Kozai-Lidov oscillations with 51 Eri b
In the Kozai-Lidov scenario, the orbit of a planet is perturbed by a distant third body, which can lead to oscillations in the inclination and eccentricity of the inner companion. The timescale of such oscillations is on the order of magnitude of (Ford and Rasio 2008):
@@
In this equation, $P_{\mathrm{Kozai}}$ is the period of the Kozai-Lidov oscillations, $M_{\star}$, the mass of the host star, $M_{\mathrm{pert}}$, $a_{\mathrm{pert}}$ and $e_{\mathrm{pert}}$ the mass, semi-major axis and eccentricity of the distant perturber, and $P_{\mathrm{planet}}$ and $a_{\mathrm{planet}}$ the period and semi-major axis of the inner planet. We computed the period of the Kozai-Lidov oscillations for different values of $M_{\mathrm{pert}}$, $a_{\mathrm{pert}}$ and $e_{\mathrm{pert}}$ using the posterior distributions of $P_{\mathrm{planet}}$ and $a_{\mathrm{planet}}$ inferred in this study. Fig. 8 shows the results of this analysis. The red curve shows the age of the system. The parameter space for Kozai-Lidov oscillations on a timescale longer than the age of the system lies above this curve. Therefore, this parameter space is not applicable. The white curve represents the SPHERE detection limits from Chomez et al. (2025). Hence, the range of possible solutions for the mass and semi-major axis of the unseen perturber lies between these 2 curves. The binary stellar companion GJ 3305 AB is overplotted. We considered the instantaneous sky-projected separation as the semi-major axis of the system and ($\approx$ 2000 AU, Feigelson et al. 2006) and the dynamical mass estimate of the binary ($\approx$ 1.1 M, Montet et al. 2015). If the binary stellar companion has a high eccentricity ($\gtrsim$ 0.9), it could potentially be responsible for the orbit of 51 Eri b, especially since, with such a high eccentricity, the semi-major axis could be significantly lower than the value of 2000 AU used in this analysis, lowering the Kozai-Lidov period. Indeed, the maximum possible separation of the stellar companion is $r_{apastron}=(1+e)*a$. This provides a lower limit for the semi-major axis of the stellar companion, i.e. $a>\frac{r_{apastron}}{1+e}$. For extreme values of the eccentricity, the minimum possible semi-major axis is $a_{min}=\frac{r_{apastron}}{2}$. This yields a range of possible values for the semi-major axis of the stellar companion, which can be as low as half of the value of the instantaneous sky-projected separation. This is a very unlikely scenario, however. Additional constrains on the possibility of GJ 3305 AB being in Kozai-Lidov resonance with 51 Eri b could be brought by a direct measurement of the RV of the stellar companion, as it would help constrain both the inclination and the longitude of ascending node of the stellar companion. The Kozai-Lidov resonance requires the mutual inclination I between the 2 objects in resonance satisfies (Kozai 1962; Lidov 1962):
@@
From the RV measurement of the stellar companion, we can better constrain its inclination and longitude of ascending node. Then, the mutual inclination between 51 Eri b and GJ 3305 AB is computed as (Murray and Dermott 2000):
@@
with i1, i2, $\Omega_{1}$, $\Omega_{2}$ the inclinations and longitude of ascending node of the 2 companions.

### v2

@@ 5 Discussion and conclusions
The implications of the planet’s high eccentricity have been extensively discussed in previous studies (Maire et al. 2019; Dupuy et al. 2021; Balmer et al. 2025). Dynamical interactions with a massive inner companion could in principle reproduce the observed eccentricity, but the required mass would exceed SPHERE detection limits (Maire et al. 2019). An alternative scenario involves secular resonances with the massive circumstellar disk, which can excite planetary eccentricities up to $\sim$0.6, even for modest disk eccentricities (Li and Lai 2023). Kozai–Lidov oscillations induced by a distant third body, such as the stellar companion GJ 3305 AB, can also generate large eccentricities (Kozai 1962; Lidov 1962). However, the wide separation of GJ 3305 AB implies a Kozai–Lidov timescale longer than the age of the system, making this scenario unlikely (Montet et al. 2015), except in the case of a very high stellar companion eccentricity ($\gtrsim 0.9$; see Appendix F). Kozai–Lidov oscillations with an undetected low-mass outer companion remain possible; given SPHERE/IRDIS detection limits (Chomez et al. 2025), such a companion would likely be Jovian or sub-Jovian ($\lesssim 1$$2,M_{\mathrm{Jup}}$). Continued astrometric and RV monitoring will further constrain this scenario (Lacour et al. 2021).
@@ Appendix B Considering whether the results of the fourth epoch are biased
To highlight these correlated structures, it is convenient to compute the autocorrelation function (ACF) of the CCF. Figure 4 presents the individual ACFs of the 4 CCFs. The first three epochs are very consistent. The last epoch, however, exhibits prominent correlated structures at velocities multiples of $\sim$ 11 $\mathrm{km}\,\mathrm{s}^{-1}$, which is consistent with the previous analysis. It also shows that the fringes cause a decrease in the full width at half-maximum (FWHM) of the ACF, compared to the other epochs. It means that the fringes bias the FWHM of the CCF to lower values. Since we expect the FWHM of the CCF and the $v\sin i$ to be correlated (see Brady et al. 2023, for example), we can conclude that the fringes likely bias the $v\sin i$ estimate of the last epoch towards lower values. However, they likely do not bias the RV as much. Provided that the template correlates better with the data than the fringes do, the CCF still provides a correct measurement of the RV of the planet.
@@
Figure 3: PSD of the data filtered from the stellar speckles estimate of our four epochs. The PSD was lowpass-filtered in order to focus on its low-frequency content. A secondary axis is displayed at the top, which corresponds to an equivalent velocity. A peak is visible around a resolution of 25 000 for each epoch, which corresponds to $\sim$12 $\mathrm{km}\,\mathrm{s}^{-1}$. This peak is particularly prominent in the data from the last epoch, when the planet was closer to the host star.
Figure 4: Normalized ACFs of the individual CCFs. The first three epochs show highly consistent ACFs, whereas the last epoch shows deeply correlated structures in the CCF at multiples of 11 km/s. The vertical dashed lines are plotted at multiples of $\pm$ 11 $\mathrm{km}\,\mathrm{s}^{-1}$ to highlight the correlated structures.
@@ Appendix C Method for computing the cross-correlation functions
To compute the cross-correlation functions, we use the formalism defined in Zucker (2003). The cross-correlation function is defined as
@@
Figure 5: Individual CCFs for the four epochs.
@@ Appendix D Orbit posterior distribution
Figure 6 presents the posterior distribution of the orbital parameters of 51 Eri b, without our RV measurements (in red) and with our RV measurements (in blue). The inclusion of the HiRISE RV measurements brings significant improvements to the orbital parameters. It also breaks the degeneracy between particular values of $\omega$ and $\Omega$ typically arising with orbital fittings using only astrometric data points.
@@ Appendix E Contrast in reflected and emitted light
with $R$ the radius of the planet, $a$ its semi-major axis, $A$ its albedo, and $f$ the phase function. We considered three different values for the phase function to show the impact of the phase of the planet. For $A$, we considered a value of 0.5, similar to that of Jupiter in visible (Karkoschka 1998). The values of $a$ and $R$ have been determined from the orbital and atmospheric fit, respectively.
@@ Appendix F Kozai-Lidov oscillations with 51 Eri b
In the Kozai-Lidov scenario, the orbit of a planet is perturbed by a distant third body, which can lead to oscillations in the inclination and eccentricity of the inner companion. The timescale of such oscillations is on the order of magnitude of (Ford and Rasio 2008)
@@
In this equation, $P_{\mathrm{Kozai}}$ is the period of the Kozai-Lidov oscillations, $M_{\star}$, the mass of the host star, $M_{\mathrm{pert}}$, $a_{\mathrm{pert}}$ and $e_{\mathrm{pert}}$ the mass, semi-major axis and eccentricity of the distant perturber, and $P_{\mathrm{planet}}$ and $a_{\mathrm{planet}}$ the period and semi-major axis of the inner planet. We computed the period of the Kozai-Lidov oscillations for different values of $M_{\mathrm{pert}}$, $a_{\mathrm{pert}}$ and $e_{\mathrm{pert}}$ using the posterior distributions of $P_{\mathrm{planet}}$ and $a_{\mathrm{planet}}$ inferred in this study. Figure 8 shows the results of this analysis. The red curve shows the age of the system. The parameter space for Kozai-Lidov oscillations on a timescale longer than the age of the system lies above this curve. Therefore, this parameter space is not applicable. The white curve represents the SPHERE detection limits from Chomez et al. (2025). Hence, the range of possible solutions for the mass and semi-major axis of the unseen perturber lies between these 2 curves. The binary stellar companion GJ 3305 AB is overplotted. We considered the instantaneous sky-projected separation as the semi-major axis of the system and ($\approx$ 2000 AU, Feigelson et al. 2006) and the dynamical mass estimate of the binary ($\approx$ 1.1 M, Montet et al. 2015). If the binary stellar companion has a high eccentricity ($\gtrsim$ 0.9), it could potentially be responsible for the orbit of 51 Eri b, especially since, with such a high eccentricity, the semi-major axis could be significantly lower than the value of 2000 AU used in this analysis, lowering the Kozai-Lidov period. Indeed, the maximum possible separation of the stellar companion is $r_{apastron}=(1+e)*a$. This provides a lower limit for the semi-major axis of the stellar companion, i.e. $a>\frac{r_{apastron}}{1+e}$. For extreme values of the eccentricity, the minimum possible semi-major axis is $a_{min}=\frac{r_{apastron}}{2}$. This yields a range of possible values for the semi-major axis of the stellar companion, which can be as low as half of the value of the instantaneous sky-projected separation. This is a very unlikely scenario, however. Additional constrains on the possibility of GJ 3305 AB being in Kozai-Lidov resonance with 51 Eri b could be brought by a direct measurement of the RV of the stellar companion, as it would help constrain both the inclination and the longitude of ascending node of the stellar companion. The Kozai-Lidov resonance requires the mutual inclination I between the 2 objects in resonance satisfies (Kozai 1962; Lidov 1962)
@@
From the RV measurement of the stellar companion, we can better constrain its inclination and longitude of ascending node. Then, the mutual inclination between 51 Eri b and GJ 3305 AB is computed as (Murray and Dermott 2000)
@@
with i1, i2, $\Omega_{1}$, $\Omega_{2}$ the inclinations and longitude of ascending node of the two companions.